# Finite-Time Analysis of Gradient Descent for Shallow Transformers

### v1

@@ 4 OVERVIEW OF MAIN RESULTS
Below we present our main results informally, which are formally established with explicit constants in Section 5.5.
@@
Assuming the target function $f^{\star}$ belongs to the sup-norm constrained Hilbert Space $\mathcal{F}_{\bar{\nu}}$ (formalized in Section 5.4) and we run ProjGD for $\tau$ steps, the average iterate $\bar{\varphi}^{(\tau)}:=\frac{1}{\tau}\sum_{s=0}^{\tau-1}\varphi^{(s)}$ satisfies
$\widehat{\mathcal{L}}_{n}\big(\bar{\varphi}$ $^{(\tau)}\big)\lesssim\underbrace{\tfrac{(d+\log(m/\delta^{\prime}))^{2}}{\sqrt{\tau}}}_{\text{optimization error}}$
$+$ $\underbrace{\sqrt{\tfrac{\log(n/\delta)}{m}}}_{\text{approximation error}}+\underbrace{\tfrac{d+\log(m/\delta^{\prime})}{\sqrt{m}}}_{\text{linearization error}}$
with an appropriate step size, projection radius, and with probability at least $1-\delta-\delta^{\prime}$ over the random initialization (introduced in Section 5.2).
For ProjSGD, the same bound holds for $\mathbb{E}[\widehat{\mathcal{L}}_{n}(\bar{\varphi}^{(\tau)})\,|\,\varphi^{(0)}]$ with the same rates (cf. Proposition 1). We highlight the following features:
•  Non-asymptotic in time and width. The training loss decays at the canonical $1/\sqrt{\tau}$ optimization rate and $1/\sqrt{m}$ width rate.
•  Logarithmic in sample size. The dependence on the sample size enters only via $\sqrt{\log(n/\delta)/m}$, i.e., logarithmically in $n$.
•  No explicit dependence on sequence length $T$. The displayed rate has no $T$ dependence thanks to the stable structure of the attention gradients.
@@ 5 TRANSFORMER NTK ANALYSIS
In this section, for the Transformer architecture introduced in Section 3.2, we analyze its gradients in Section 5.1, introduce the initialization in Section 5.2, derive the Transformer NTK in Section 5.3, and introduce the concept of transportation mappings in Section 5.4. These tools define the function class against which our main results are established in Section 5.5. Proofs of the stated results are provided in the appendices as part of the supplementary material.
@@ 5.1 Transformer Gradient Analysis
With $\alpha=\sigma_{s}(X^{\top}W_{i}q_{X})\in\Delta_{T}$ and $\mu=\sum_{t}\alpha_{t}X_{t}$, notice that $M(X;W_{i})=\sum_{t=1}^{T}\alpha_{t}X_{t}X_{t}^{\top}-\mu\mu^{\top}$ is a softmax‑weighted covariance that appears in (7) as a result of the softmax Jacobian matrix. Since $\mathrm{tr}\,M(X;W_{i})\leq 1$ for all $X\in\mathcal{X}$, we obtain $\|M(X;W_{i})\|_{F}\leq 1$ uniformly in $T$. As a consequence of this key property, we observe that $\|\nabla_{W_{i}}f(X;\varphi)\|_{F}$ is independent of $T$. In contrast, for RNNs the analogous gradient contains an explicit sum over $T$ recurrent steps, leading to constants that can grow with $T$; see (Cayci and Eryilmaz, 2024a, Prop. 3.1). This structural difference underlies our TT‑independent optimization bounds in Section 5.5.
@@ 5.2 Initialization
By Gaussian concentration for the $1$-Lipschitz map $u\mapsto\|u\|_{2}$ (see (Boucheron et al., 2013, Theorem 5.6)) and a union bound over $i\in[m/2]$, we obtain $\mathbb{P}\big(\mathcal{E}_{U,m}(\delta^{\prime})\big)\;\geq\;1-\delta^{\prime}.$
@@
$\Omega_{\rho,i}\;:=\;\mathcal{C}_{\rho_{c},i}\times\mathcal{U}_{\rho_{u},i}\times\mathcal{W}_{\rho_{w},i},\qquad\Omega_{\rho}\;:=\;\prod_{i=1}^{m}\Omega_{\rho,i}.$
Remark 7.
@@
$\|U_{i}\|_{2}\leq\sqrt{d}+\sqrt{2\log(m/2\delta^{\prime})}+\frac{\rho_{u}}{\sqrt{m}}=:B_{U,m}(\delta^{\prime}).$
This allows us to establish the local Lipschitz continuity and smoothness, which in turn enable us to control the optimization error in Theorem 1.
@@
Next, we define the linear model, which will be useful in the analysis.
@@ 5.3 Transformer NTK
Let $\varphi^{(0)}$ be the random initialization. The NTK is defined as
@@
Because $\nabla_{\varphi}f(X;\varphi^{(0)})$ is block-separable across $(c,U,W)$, we investigate the features coming from these parameters separately. Using the gradient identities in Lemma 1, the Transformer NTK decomposes into the following blocks.
@@ 5.4 Transportation Mappings
Define a measurable $v:\mathbb{R}^{d^{2}+d+1}\to\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ with $v(\varphi_{0})=(v_{c}(\varphi_{0}),v_{u}(\varphi_{0}),v_{w}(\varphi_{0}))$ such that $\mathbb{E}\!\Big[\,|v_{c}(\varphi_{0})|^{2}+\|v_{u}(\varphi_{0})\|_{2}^{2}+\|v_{w}(\varphi_{0})\|_{F}^{2}\,\Big]<\infty$. Define the function class
@@
Note that we have $\|\tilde{f}\|_{\mathcal{F}_{\bar{\nu}}}\leq\|\bar{\nu}\|_{2}$. This constraint enables us to control the approximation bound in Lemma 5 and gives a significant simplification in the convergence analysis of ProjGD in Theorem 1.
@@
$\varepsilon_{\mathrm{app}}\;\leq\;2(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{2\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$
@@
Contrasting this result with the RNN counterpart (Cayci and Eryilmaz, 2024a, Proposition 3.5), we can see that stable attention gradient norms give an approximation bound that does not depend on $T$.
@@ 5.5 Main Results
Under the same conditions as Theorem 1, we have
@@ 6 EXPERIMENTAL RESULTS
In Section 6.1, we validate the scaling laws predicted by Lemma 3, Lemma 5, and Theorem 1. In Section 6.2, we compare IndRNN and our Transformer model on an autoregressive forecasting task, highlighting differences in handling long-term dependencies and the accompanying memory trade-offs.
@@ 6.1 Teacher–Student Scaling-Law Validation
We assess the three components of our theory—linearization, approximation, and optimization—with controlled teacher–student experiments using the same shallow Transformer as in Section 3.2 with $q_{X}=X_{T}$. Throughout, inputs are i.i.d. $X\sim\mathcal{N}(0,I)$ with token-wise clipping to keep $\max_{t}\|X_{t}\|_{2}\leq 1$.
@@
$\tilde{y}^{(j)}=\frac{1}{\tilde{m}}\sum_{i=1}^{\tilde{m}}\!\big[\phi_{c}(X^{(j)})\,v_{c}(\varphi_{i}^{(0)})+\langle\phi_{u}(X^{(j)}),v_{u}(\varphi_{i}^{(0)})\rangle\\ +\langle\phi_{w}(X^{(j)}),v_{w}(\varphi_{i}^{(0)})\rangle_{F}\big],$
@@
Students are initialized with the symmetric initialization from Section 5.2 so that $f(\cdot;\varphi^{(0)})\equiv 0$. For each width $m\in\{8,16,64,128,256\}$, we record:
@@
We average curves over $10$ seeds, show $\pm$95% confidence interval (CI) around the mean, and fit a straight line to the log–log curves to estimate the scaling exponent. It can be seen from Figure 1 that all three quantities—linearization error, approximation error, and minimum training loss—decrease monotonically with width and exhibit a clear power-law dependence on $m$. The fitted slopes are in close agreement with the predicted $m^{-1/2}$ rate. Medians closely track the mean and the CIs remain small across widths, indicating that the observed scaling is robust to random initialization.
@@
To compare how a recursive model (IndRNN (Cayci and Eryilmaz, 2024a)) and our Transformer cope with long-term dependencies and training stability, we use a controlled autoregressive AR($L$) task where the lag $L$ sets the long-term dependency of the problem. Theory for diagonal RNNs predicts that depending on a Lipschitz cutoff, gradients can grow exponentially due to repeated multiplication over a horizon of $T$ steps (Cayci and Eryilmaz, 2024a), whereas our Transformer’s bound has no explicit $T$-dependence (cf. Lemma 2). Each input sequence $X\!\in\!\mathbb{R}^{1\times T}$ is constructed as a window from a univariate process
@@
and the label is $y=X_{T+1}$. To the input we append fixed positional channels ($d_{\text{pos}}{=}8$ sinusoidal pairs) so that the Transformer model has positional knowledge. We fix n=5,000n=5{,}000, $\alpha=0.9$, $\sigma_{\varepsilon}^{2}=0.1$, and vary the lag $L$ to toggle short/long temporal dependency inside the window. For Transformer, we set $T=L+1$ so that the lag is inside the context window, which makes Transformer’s memory footprint $\mathcal{O}(L)$ compared to IndRNN’s $\mathcal{O}(1)$ footprint. Both models are initialized with a symmetric initialization and are trained with gradient descent for $\tau=2{,}000$ steps with step size $\eta=1/\sqrt{\tau}$ and no projection to observe the gradient norm dynamics. We set $m=64$ and use $\tanh$. The results are averaged over $20$ seeds.
Figure 2: Performance comparison of IndRNN and Transformer architectures with varying time lag $L$.
In Figure 2, we contrast the function Jacobian norms and the minimum validation loss performances of the IndRNN and Transformer architectures for a range of lag values. In particular, the upper plot compares the maximum Jacobian norm occurred in training with respect to the diagonal recurrent weights for IndRNN and attention weights for Transformer. We observe that maximum IndRNN norms are large with higher fluctuations, which confirms the theory established in (Cayci and Eryilmaz, 2024a); whereas Transformer norms consistently take smaller values as predicted by our theory. The lower plot compares the minimum validation loss occurred in training for both architectures. It can be seen that when the problem has a short-term dependency (low $L$), IndRNN achieves better results with degrading performance as $L$ increases; on the other hand, Transformer gives a consistent performance that does not degrade with $L$, at the expense of its growing memory footprint. As such, this highlights the performance-memory trade-off between the two architectures that is anticipated by the theory.

### v2

@@ 4 OVERVIEW OF MAIN RESULTS
Below we present our main results informally. Formal statements are given in Section 5.5, with detailed proofs deferred to Appendix A.
@@
Let $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ be the target function (formalized in Section 5.4). After $\tau$ steps of ProjGD with an appropriate step size and projection radius, the average iterate $\bar{\varphi}^{(\tau)}:=\frac{1}{\tau}\sum_{s=0}^{\tau-1}\varphi^{(s)}$ satisfies, with probability at least $1-\delta-\delta^{\prime}$ over the random initialization (introduced in Section 5.2),
$\widehat{\mathcal{L}}_{n}\!\left(\bar{\varphi}^{(\tau)}\right)\lesssim\underbrace{\frac{D^{2}}{\sqrt{\tau}}}_{\begin{subarray}{c}\text{optimization}\\ \text{error}\end{subarray}}+\underbrace{\sqrt{\frac{D\log(n/\delta)}{m}}}_{\begin{subarray}{c}\text{approximation}\\ \text{error}\end{subarray}}+\underbrace{\sqrt{\frac{D^{3}}{m}}}_{\begin{subarray}{c}\text{linearization}\\ \text{error}\end{subarray}}$
where $D:=d+\log(m/\delta^{\prime})$.
For ProjSGD, the same bound holds for $\mathbb{E}[\widehat{\mathcal{L}}_{n}(\bar{\varphi}^{(\tau)})\,|\,\varphi^{(0)}]$ with the same rates (cf. Proposition 1).
We highlight the following features:
•  Non-asymptotic in time and width. The training loss admits a finite-time optimization–approximation–linearization decomposition, with the canonical $1/\sqrt{\tau}$ optimization rate and $1/\sqrt{m}$ width dependence. This decomposition holds for any $\tau,m\geq 1$, with no hidden asymptotic regime.
•  Logarithmic in sample size. The sample size enters only through the approximation error (Lemma 5). To make this term at most $\varepsilon$ it suffices to take $m\gtrsim\varepsilon^{-2}\log(n/\delta)$; in particular, $m$ needs to grow only logarithmically with $n$.
•  No explicit dependence on sequence length. Thanks to the attention gradients being uniformly bounded in $T$ via a softmax-covariance structure (Remark 3), we obtain an optimization error independent of the sequence length $T$.
@@ 5 TRANSFORMER NTK ANALYSIS
In this section, we analyze the gradients of the Transformer architecture from Section 3.2, specify the initialization, derive the associated NTK, and introduce transportation mappings. These components define the function class used to establish our main results in Section 5.5.
@@ 5.1 Transformer Gradient Analysis
With $\alpha=\sigma_{s}(X^{\top}W_{i}q_{X})\in\Delta_{T}$ and $\mu=\sum_{t}\alpha_{t}X_{t}$, notice that $M(X;W_{i})=\sum_{t=1}^{T}\alpha_{t}X_{t}X_{t}^{\top}-\mu\mu^{\top}$ is a softmax‑weighted covariance matrix. This is the only term in Lemma 1 that involves a $T$-dependence. Since $\mathrm{tr}\,M(X;W_{i})\leq 1$ for all $X\in\mathcal{X}$, we obtain $\|M(X;W_{i})\|_{F}\leq 1$ uniformly in $T$. As a consequence of this key property, we observe that $\|\nabla_{W_{i}}f(X;\varphi)\|_{F}$ is independent of $T$. In contrast, for RNNs the analogous gradient contains an explicit sum over $T$ recurrent steps, leading to constants that can grow with $T$ (Cayci and Eryilmaz, 2024a, Prop. 3.1). This structural difference underlies our TT‑independent optimization bounds in Section 5.5. See Appendix A.1 for the detailed derivation.
@@ 5.2 Initialization
By Gaussian concentration for the $1$-Lipschitz map $u\mapsto\|u\|_{2}$ (Boucheron et al., 2013, Thm. 5.6) and a union bound over $i\in[m/2]$, we obtain $\mathbb{P}\big(\mathcal{E}_{U,m}(\delta^{\prime})\big)\;\geq\;1-\delta^{\prime}.$
@@
$\Omega_{\rho,i}\,:=\,\mathcal{C}_{\rho_{c},i}\times\mathcal{U}_{\rho_{u},i}\times\mathcal{W}_{\rho_{w},i},\quad\,\Omega_{\rho}\,:=\,\prod_{i=1}^{m}\Omega_{\rho,i}.$
@@
$\|U_{i}\|_{2}\leq\sqrt{d}+\sqrt{2\log(m/2\delta^{\prime})}+\frac{\rho_{u}}{\sqrt{m}}=:B_{U,m}(\delta^{\prime}).$ (9)
This allows us to establish local Lipschitz continuity and smoothness, as formalized in the following lemma, which in turn enables control of the optimization error in Theorem 1.
@@
Next, we define the linear model, which will be useful in the analysis to characterize the linearization error.
@@ 5.3 Transformer NTK
For our convergence analysis, we work in the NTK regime, where training is well-approximated by linearization around initialization. Concretely, for a given random initialization $\varphi^{(0)}$, the Neural Tangent Kernel (NTK) is defined as
@@
Since $\nabla_{\varphi}f(X;\varphi^{(0)})$ is block-separable across $(c,U,W)$, we investigate the features coming from these parameters separately. Using the gradient identities in Lemma 1, the Transformer NTK decomposes into the following blocks.
@@ 5.4 Transportation Mappings
Let $v:\mathbb{R}^{d^{2}+d+1}\to\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ be a measurable function with $v(\varphi_{0})=(v_{c}(\varphi_{0}),v_{u}(\varphi_{0}),v_{w}(\varphi_{0}))$ such that $\mathbb{E}\!\Big[\,|v_{c}(\varphi_{0})|^{2}+\|v_{u}(\varphi_{0})\|_{2}^{2}+\|v_{w}(\varphi_{0})\|_{F}^{2}\,\Big]<\infty$. Define the function class
@@
Note that for any $\tilde{f}\in\mathcal{F}_{\bar{\nu}}$, we have $\|\tilde{f}\|_{\mathcal{F}_{\bar{\nu}}}\leq\|\bar{\nu}\|_{2}$. Thus, $\|\bar{\nu}\|_{2}$ quantifies the complexity of the function class $\mathcal{F}_{\bar{\nu}}$ and serves as a natural measure of the hardness of the regression problem. It will play a central role in controlling the approximation error (Lemma 5) and the convergence of ProjGD (Theorem 1).
@@
$\varepsilon_{\mathrm{app}}\leq 4(\sigma_{0}\bar{\nu}_{c}+\sigma_{1}\bar{\nu}_{u}+\sigma_{1}\bar{\nu}_{w})\sqrt{\tfrac{\log(2n/\delta)}{m}}=:B_{\mathrm{app},m}(\delta),$
@@
In contrast to the RNN counterpart (Cayci and Eryilmaz, 2024a, Proposition 3.5), where gradient norms may grow with the sequence length, the softmax-covariance structure of attention yields an approximation bound independent of $T$.
@@ 5.5 Main Results
$\widehat{\mathcal{L}}_{n}\!\big(\bar{\varphi}^{(\tau)}\big)\;\lesssim_{\bar{\sigma},\bar{\nu},\rho}\!\;\frac{L_{1,m}(\delta^{\prime})^{4}}{\sqrt{\tau}}+L_{1,m}(\delta^{\prime})\big(B_{\mathrm{app},m}(\delta)\\ +B_{\mathrm{lin},m}(\delta^{\prime})+\varepsilon_{\mathrm{CoF}}\big)+B_{\mathrm{lin},m}(\delta^{\prime})^{2},$
@@
Remark 9.
The realizability condition $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ is used only in the approximation step (Lemma 5) to control how well $f^{\star}$ can be represented in the Transformer-NTK RKHS. If $f^{\star}$ is not realizable by $\mathcal{F}_{\bar{\nu}}$, the optimization and linearization arguments (Lemmas 23) and the Lyapunov analysis in Theorem 1 go through unchanged, and the final bound acquires the additional term
$\inf_{\tilde{f}\in\mathcal{F}_{\bar{\nu}}}\ \max_{j\in[n]}\big|f^{\star}(X^{(j)})-\tilde{f}(X^{(j)})\big|.$
We leave a detailed study of the approximation of broader target functions by the Transformer-NTK RKHS for future work.
Remark 10.
The condition $\rho\geq\bar{\nu}$ ensures that the transported parameters $\tilde{\varphi}$ constructed from $v\in\mathcal{V}_{\bar{\nu}}$ (cf. (11)) lie in the feasible set $\Omega_{\rho}$. Larger $\bar{\nu}$ corresponds to a richer target class (larger RKHS norm budget cf. Remark 7) and requires a larger projection radius, which in turn increases the hidden constants in Theorem 1. See Appendix A.7 for explicit dependencies.
@@
Running ProjSGD under the same setup as Theorem 1 gives, on the event $\mathcal{E}_{U,m}(\delta^{\prime})\cap\mathcal{E}_{\mathrm{app}}(\delta)$,
@@
Comparison of results.
Theorem 1 gives a high-probability finite-time bound for ProjGD, stated for the minimum iterate and the averaged iterate. Proposition 1 shows that ProjSGD enjoys the same error decomposition and the same rates in $(\tau,m,n)$, with the guarantee stated in conditional expectation over the sampling randomness.
@@ 6 EXPERIMENTAL RESULTS
We complement the theory with two experiments. Section 6.1 validates the scaling laws predicted by Lemma 3, Lemma 5, and Theorem 1 in a realizable teacher–student setting aligned with $f^{\star}\in\mathcal{F}_{\bar{\nu}}$ and near-initialization assumptions. Section 6.2 then probes behavior beyond this regime by comparing IndRNNs and our Transformer on an autoregressive AR($L$) task trained without projection, highlighting long-term dependency behavior and the associated memory trade-off.
@@ 6.1 Teacher–Student Scaling-Law Validation
We assess the three components of our theory—linearization, approximation, and optimization—with controlled teacher–student experiments using the same shallow Transformer as in Section 3.2 with $q_{X}=X_{T}$. Throughout, inputs are i.i.d. $X_{t}\sim\mathcal{N}(0,I_{d})$ with token-wise clipping to keep $\max_{t}\|X_{t}\|_{2}\leq 1$.
@@
$\tilde{y}^{(j)}=\frac{1}{\tilde{m}}\sum_{i=1}^{\tilde{m}}\!\big[\phi_{c}(X^{(j)})\,v_{c}(\varphi_{i}^{(0)})\\ +\langle\phi_{u}(X^{(j)}),v_{u}(\varphi_{i}^{(0)})\rangle+\langle\phi_{w}(X^{(j)}),v_{w}(\varphi_{i}^{(0)})\rangle_{F}\big],$
@@
Students are initialized with the symmetric initialization from Section 5.2 so that $f(\cdot;\varphi^{(0)})\equiv 0$. For each width $m\in\{8,16,32,64,128,256\}$, we record:
@@
We average curves over $10$ seeds, show $\pm$95% confidence interval (CI) around the mean, and fit a straight line to the log–log curves to estimate the scaling exponent. Figure 1 shows that all three quantities—linearization error, approximation error, and minimum training loss—decrease monotonically with width and exhibit a clear power-law dependence on $m$. The fitted slopes are in close agreement with the predicted $m^{-1/2}$ rate. Medians closely track the mean and the CIs remain small across widths, indicating that the observed scaling is robust to random initialization.
@@
To compare how IndRNNs (Cayci and Eryilmaz, 2024a) and our Transformer cope with long-term dependencies and training stability, we use a controlled autoregressive AR($L$) task where the lag $L$ sets the dependency horizon. Theory for diagonal RNNs predicts that, depending on a Lipschitz cutoff, gradients can grow exponentially due to repeated multiplication over a horizon of $T$ steps (Cayci and Eryilmaz, 2024a), whereas our Transformer’s bound has no explicit $T$-dependence (cf. Lemma 2).
Each input sequence $X\!\in\!\mathbb{R}^{1\times T}$ is constructed as a window from a univariate process
@@
and the label is $y=X_{T+1}$. We append fixed positional channels ($d_{\text{pos}}{=}8$ sinusoidal pairs) to the input so that the Transformer has positional information. We fix n=5,000n=5{,}000, $\alpha=0.9$, $\sigma_{\varepsilon}^{2}=0.1$, and vary $L$. For the Transformer we set $T=L+1$ so that the lag falls within the context window, yielding a memory footprint $\mathcal{O}(L)$ compared to IndRNN’s $\mathcal{O}(1)$. Both models use symmetric initialization and are trained with gradient descent for $\tau=2{,}000$ steps with step size $\eta=1/\sqrt{\tau}$ and no projection to observe gradient-norm dynamics. We set $m=64$ and use $\tanh$ as the activation. Results are averaged over $20$ seeds.
Metrics.
We track training stability via the maximum Jacobian norm over training, computed with respect to the diagonal recurrent weights (IndRNN) or the attention weights (Transformer). Predictive performance is summarized by the minimum validation loss achieved over training.
Figure 2 summarizes the comparison across lags $L$. The top panel shows that IndRNN exhibits larger and more variable norms, consistent with the potential gradient growth predicted in Cayci and Eryilmaz (2024a), whereas the Transformer remains stable. The bottom panel shows the best validation loss achieved over training: IndRNN performs well for small lags but degrades as $L$ increases, while the Transformer maintains roughly constant performance as $L$ grows, at the cost of an $\mathcal{O}(L)$ context window. Overall, the experiment illustrates the performance-memory trade-off between the two architectures anticipated by the theory. Although this task lies outside our realizable/projection regime, the observed stability of attention is consistent with the softmax-covariance control highlighted in Remark 3.
Figure 2: IndRNN vs. Transformer on the AR($L$) task with $T=L+1$. Top: maximum Jacobian norm over training (stability); bottom: minimum validation loss achieved over training (performance). Results are averaged over $20$ seeds.